consider c=200+0.8(Y-T)
what is the national income model ?
what is the equilibrium level of national income
1
Expert's answer
2014-09-15T13:20:33-0400
The national-income model is Y = C + I + G
The equilibrium level of national income will be: C = 200 + 0.8(Y - T), so 0.8Y = C + 0.8T - 200, Y = 1.25C + T - 250